OpenStudy (anonymous):

Find the slope of the line that passes through the origin and (8, -5).

OpenStudy (anonymous):

The origin is the point (0,0)

hero (hero):

y = mx + b 0 = 0m + b -5 = 8m + b b = 0 -5 = 8m + 0 -5 = 8m -5/8 = m